write an equation of the line containing the given point and parallel to the given line, express your answer in the form y=mx+b
(-5,7); 2x=9y+5
Thank you

Answer by stanbon(75887)

write an equation of the line containing the given point and parallel to the given line, express your answer in the form y=mx+b
(-5,7); 2x=9y+5
-------------
If you put the given line in slope-intercept form
you will see its slope is (2/9)
----
You want the equation of a line with slope
= 2/9 where y=7 when x = -5.
---
Form:
y = mx+b
7 = (2/9)(-5)+b
b = 63/9 +10/9
b = 73/9
----------------
y = (9/2)x + (73/9)
====
Cheers,
stan H.